OpenStudy (anonymous):

A hot air balloon holds 2,200 cubic meters of helium. The density of helium is 0.1785 kilograms per cubic meter. How many kilograms of helium does the balloon contain, rounded to the nearest tenth of a kilogram? 392.7 kg 2,199.8 kg 2,200.2 kg 11,195.9 kg

OpenStudy (anonymous):

@paki

OpenStudy (paki):

use formula.... density=mass/volume

OpenStudy (paki):

@shinebrightlikeadimon

OpenStudy (anonymous):

but i don't know which numbers to plug in and where to plug in

OpenStudy (paki):

what is density here.... what is mass here... what is volume her...

OpenStudy (paki):

read the question...and tell me...

OpenStudy (paki):

@shinebrightlikeadimon

OpenStudy (paki):

ok see.. here: density is 0.1785 kilograms per cubic meter volume is 2,200 cubic meters and we need mass... so just put all these in the formula and solve... @shinebrightlikeadimon

OpenStudy (anonymous):

k give me a second

OpenStudy (paki):

yeah do it...

OpenStudy (anonymous):

392.7? is what i got

OpenStudy (anonymous):

cause i multiplied them togeth i did 0.1785= mass/2200 and i had to find m so i did 0.1785*2200 = mass/ 2200*2200 392.7 = mass

OpenStudy (paki):

yeah rite...

OpenStudy (paki):

good

OpenStudy (anonymous):

so A?

OpenStudy (paki):

yes

OpenStudy (anonymous):

i need help with more can you help?

OpenStudy (paki):

yeah... i will try... and close this question too...